OpenStudy (anonymous):

If the price of a magazine is 60% higher than the price of a newspaper, then the price of the newspaper is : A. 37.5% lower than the price of the magazine B. 40% lower than the price of the magazine C. 60% lower than the price of the magazine D. 62.5% lower than the price of the magazine

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (mayankdevnani):

let the price of newspaper be x and magazine be y. so, \[\huge \bf x+60~percent=y\]

OpenStudy (mayankdevnani):

Can you solve for x ? @kryton1212

OpenStudy (anonymous):

If \(n\) is the price of the newspaper and \(m\) is the price of the magazine, then...

OpenStudy (anonymous):

\[ m = n + 60\% n\]Which simplifies to: \[m = 1.6 n \]We want to find the number such that: \[ m - ??\%m = n \]Well, that is my interpretation.

OpenStudy (anonymous):

Maybe I'm reading into it too much.

OpenStudy (anonymous):

One thing that would make it easy is to simply say that the newspaper is priced at $1.00

OpenStudy (anonymous):

Thus the magazine is $1.60

OpenStudy (anonymous):

Then we have $1.60 - $1.60x = $1.00

OpenStudy (mayankdevnani):

@wio you go into deep .... the answer is easy. its option C) let the price of newspaper be x and magazine be y, so, \[\huge \bf x+60~percent=y\] subtract 60 % from both the sides, we get \[\huge \bf x=y-60~percent\] and x=newspaper and y=magazine.

OpenStudy (mayankdevnani):

its too simple. @kryton1212 and @wio

OpenStudy (anonymous):

60% of what though?

OpenStudy (mayankdevnani):

its 60% higher than newspaper

OpenStudy (mayankdevnani):

and its price.

OpenStudy (mayankdevnani):

@wio understood?

OpenStudy (anonymous):

\[ $1.60 - $1.60x = $1.00\\ $0.60 = $1.60x \\ 0.375 = x \]So I get 37.5%

OpenStudy (mayankdevnani):

Why shouldn't it B? @wio

OpenStudy (anonymous):

You're saying it is 60% of \(x\), and if so, then your formula should be \[ x + 60\%x = y \]Because percentages are multiplied to something.

OpenStudy (mayankdevnani):

its 60% higher.... it is obvious added

OpenStudy (anonymous):

So the question is asking \[ x = y - ??\%y \]

OpenStudy (anonymous):

If an apple cost a dollar and a banana is 60% higher... what is the price of the banana?

OpenStudy (mandre):

m = 1.6 * n m = n / 1.6 \[n = \frac{ m }{ 1.6 }\] \[n = \frac{ m }{ 16/10 }\] \[n = m(10/16)\] \[n = 0.625m\] To me that means the newspaper is (1-0.625) = 37.5% lower than the magazine.